OpenStudy (anonymous):

How do I simplify this exponent? [(−6x2y)3]

OpenStudy (anonymous):

\[[(-6x^{2}y)^{3}]\]

OpenStudy (anonymous):

just multiply it out... \(\ (-6x^2y)^3= (-6x^2y)(-6x^2y)(-6x^2y)=(-6 \cdot -6 \cdot -6)(x^2 \cdot x^2 \cdot x^2)(y \cdot y \cdot y) \) = ????

OpenStudy (anonymous):

Thanks. I think I need to talk to my teacher. I still don't get it.